OpenStudy (anonymous):

which of the following points lie in the solution set to the following system of inequalities y<3x+3, y

OpenStudy (anonymous):

here, its better to verify the options to get answer quickly. take the 1st option (1,-5) here x=1 and y=-5 3x+3=3(1)+3=6 > -5 => -5<6 1+2=3 >-5 => -5<3 hence both conditions r satisfied

OpenStudy (precal):



OpenStudy (precal):

test (0,0) for both 0<3 true shade down 0<2 shade down b/c is it true

OpenStudy (precal):